Lindiwe de



ides that she would like to buy her daughter, Mbali, a



ar when she turns 21 in six years' time.



She deposits R6 000 ea



h month into an a




ount earning 8,94% interest per year,



ompounded monthly. The



amount that Lindiwe (rounded to the nearest rand) will have available six years from now is




1
Expert's answer
2022-03-15T18:08:54-0400

it's an annuity


"FV=6000[(1+\frac{0.0894}{12})^{12×6}-1÷0.00745]"


"FV=6000(\frac{(1+0.00745)^{72}-1}{0.00745})=R568948.33"